OpenStudy (coles):

What is -√180 simplifyed

OpenStudy (anonymous):

13.4164079 i

OpenStudy (anonymous):

If the negative sign is supposed to be inside the radical, then you can break it down like so:\[\sqrt{-180} = i \sqrt{180} = 2i \sqrt{45} = 6i \sqrt{5}\]

OpenStudy (anonymous):

@Krissy1997, that's not simplified, it's estimated.